OpenStudy (anonymous):

What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (anonymous):

first you need the equation of the line the slope is \[m=\frac{y_2-y_1}{x_2-x_1}=\frac{3+3}{-5-1}=\frac{6}{-6}=-1\]

OpenStudy (anonymous):

then you can use the point slope formula \(y-y_1=m(x-x_1)\) with \(m=-1, x_1=-5,y_1=3\)